OpenStudy (anonymous):

hi guys...i want to ask questions about integration ...it's just a basic but i really2x 4gt how to solve it...here's the question. Integration of 2z-4 over 5z-9 dz help pls...

OpenStudy (raden):

integral by subs... let u = 5z-9 --------> z=(u+9)/5 du = 5 dz dz = 1/5 du so, it can be : int (2(u+9)/5 - 4))/u * 1/5 du simplify it ....

OpenStudy (anonymous):

thank you...let me try first...

OpenStudy (anonymous):

i got it ...thank you very much raden !!!

OpenStudy (anonymous):

i learned it last time when i was i n matric but i forgot now...hahaha

OpenStudy (raden):

well, after that the integration can be